英語 (English)

発信型英語 (Communicative English) (総科 ((社創 A)2 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 2 単位前期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This Communicative English class planned to give a good assistance

in rising up English speaking and communicating abilities. 【授業の概要】Classes will be running in ENGLISH language only for more

practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】Students will develop confidence in oral communication and refine

their listening skills. 【授業の計画】

1. Overall Guidance

2. Conversational Topic 1

3. Conversational Topic 2

4. Conversational Topic 3

5. Conversational Topic 4

6. Conversational Topic 5

7. Conversational Topic 6

8. Conversational Topic 7

9. Conversational Topic 8

10. Conversational Topic 9

11. Conversational Topic 10

12. Conversational Topic 11

13. Conversational Topic 13

14. Conversational Topic 14

15. Review 【教科書】There is no specific text book but there will be some other media

depending on the ’Topics’ which will be selected up to date. 【参考書等】 [参考資料] 【成績評価の方法】 Attendance, Homework fulfillment and Participation in class. 【再試験の有無】 No make-up examination for this course. 【受講者へのメッセージ】 Punctuality and regular attendance are extremely important. Up to three absences will be permitted. Students are expected to communicate actively with all members of the class in English.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220886 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 座喜 . (オフィスアワー: contact E-mail: jzaki@ias.tokushima-u.ac.jp)

English

Communicative English (総科 ((社創 A)2 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 2 units 前期火 7 ・ 8

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This Communicative English class planned to give a good assistance in

rising up English speaking and communicating abilities. Outline⟩ Classes will be running in ENGLISH language only for more practices

chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students will develop confidence in oral communication and refine their

listening skills.

Schedule⟩

1. Overall Guidance

2. Conversational Topic 1

3. Conversational Topic 2

4. Conversational Topic 3

5. Conversational Topic 4

6. Conversational Topic 5

7. Conversational Topic 6

8. Conversational Topic 7

9. Conversational Topic 8

10. Conversational Topic 9

11. Conversational Topic 10

12. Conversational Topic 11

13. Conversational Topic 13

14. Conversational Topic 14

15. Review Textbook⟩ There is no specific text book but there will be some other media

depending on the ’Topics’ which will be selected up to date. Reference⟩ [参考資料] Evaluation Criteria⟩ Attendance, Homework fulfillment and Participation in class. Re-evaluation⟩ No make-up examination for this course. Message⟩ Punctuality and regular attendance are extremely important. Up to three absences will be permitted. Students are expected to communicate actively with all members of the class in English.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220886 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki . (Office Hour: contact E-mail: jzaki@ias.tokushima-u.ac.jp)